Skip to main contentSkip to solution

A and B are two independent events. The probability that both events A and B occur is 16\frac{1}{6} and the probability that neither of them occur is 13\frac{1}{3}. If P(A) = x, P(B) = y then the value of x+y is.

Solution

✅ Correct Option: 2

A and B are independent events. The probability that both events A and B occur is 16\frac{1}{6} and the probability that neither of them occur is 13\frac{1}{3}. Given P(A) = x and P(B) = y.

Since A and B are independent:

P(A and B)=P(A)×P(B)P(A \text{ and } B) = P(A) \times P(B)

xy=16xy = \frac{1}{6}


For the probability that neither occurs, since A and B are independent, A' and B' are also independent:

P(A′ and B′)=P(A′)×P(B′)P(A' \text{ and } B') = P(A') \times P(B')

Where P(A′)=1−xP(A') = 1 - x and P(B′)=1−yP(B') = 1 - y:

(1−x)(1−y)=13(1-x)(1-y) = \frac{1}{3}


Expanding the equation:

(1−x)(1−y)=13(1-x)(1-y) = \frac{1}{3}

1−x−y+xy=131 - x - y + xy = \frac{1}{3}


Substituting xy=16xy = \frac{1}{6}:

1−x−y+16=131 - x - y + \frac{1}{6} = \frac{1}{3}

76−(x+y)=13\frac{7}{6} - (x+y) = \frac{1}{3}

76−(x+y)=26\frac{7}{6} - (x+y) = \frac{2}{6}

x+y=76−26x + y = \frac{7}{6} - \frac{2}{6}

x+y=56x + y = \frac{5}{6}

Therefore, the value of x + y is 56\frac{5}{6}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question